OpenStudy (snitchseeker1496):

Find the solution: ysec^(2) xdx - cot^(3) xln^(2)ydy=0 Y(pi/4) = e

OpenStudy (alekos):

have you tried this?

OpenStudy (sshayer):

\[y \sec ^2x~dx=\cot ^3x~\ln^2y~dy\] divide by \[y \cot ^3x\] \[\tan ^3x \sec ^2x~dx=\frac{ \ln ^2y~dy }{ y }\] check the statement?

OpenStudy (sshayer):

is it ln y^2?

OpenStudy (holsteremission):

Either way, a substitution of \(\ln y\) makes integration of the RHS easy.

OpenStudy (sshayer):

correct.